# Market Sensitivities and Growth Differentials Across Australian Housing Markets

### v2

@@
December 3, 2025
Australian house prices have risen strongly since the mid-1990s, but growth has been highly uneven across regions. Raw growth figures obscure whether these differences reflect persistent structural trends or cyclical fluctuations. We address this by estimating a three-factor model in levels for regional repeat-sales log price indexes over 1995–2024. The model decomposes each regional index into a national Market factor, two stationary spreads (Mining and Lifestyle) that capture mean-reverting geographic cycles, and a city-specific residual. The Mining spread, proxied by a Perth–Sydney index differential, reflects resource-driven oscillations in relative performance; the Lifestyle spread captures amenity-driven coastal and regional cycles. The Market loading $\beta_{r}$ isolates each region’s fundamental sensitivity to national growth, so that an assumed national change $f_{M}$ implies an individual city’s growth of $f_{r}\approx f_{M}^{\beta_{r}}$ once mean-reverting spreads are netted out. Comparing realised paths to these factor-implied trajectories indicates when a city is historically elevated or depressed, and attributes the gap to Mining or Lifestyle spreads.
Expanding-window ARIMAX estimation reveals that Market betas are stable across major shocks (the mining boom, the Global Financial Crisis, and COVID-19), while Mining and Lifestyle behave as stationary spreads that widen forecast funnels without overturning the cross-sectional ranking implied by $\beta_{r}$. Melbourne amplifies national growth, Sydney tracks the national trend closely, and regional areas dampen it. The framework thus provides a simple, factor-based tool for interpreting regional growth differentials and their persistence.
@@ 1 Introduction
Local housing markets are generally shaped by region-specific conditions, while sharing a common national price trend. Price trajectories can diverge substantially, as evidenced by asynchronous booms across U.S. cities during the 2000s housing cycle (ferreira_gyourko2011). More broadly, some of these metropolitan areas have sustained house price growth rates above the national average over very long periods of time (gyourko_mayer_sinai2013). Similarly, Australian house price appreciation has experienced considerable variation across the capital cities over decades.
This paper studies such divergence for Australian houses (excluding units and apartments) over 1995–2024, using nominal repeat-sales log price indexes. We examine how much of the long-run divergence across cities reflects broadly shared but geographically varying cyclical forces, and how much instead reflects each city’s long-run sensitivity to the national market trend. These regional growth differentials matter for housing policy, investment decisions, and spatial inequality: persistent gaps can impede internal migration and exacerbate local skill shortages, while reliable planning requires understanding whether relative growth relationships are stable.
Our approach separates a trend correlated to the national price index from two economically interpretable, mean-reverting spreads. The Mining spread captures the geographic variation in exposure to the resource cycle that raises prices more in mining-exposed regions, relative to the national market and less in Sydney, consistent with its negative Mining factor exposure in both our factor model and the PCA of sijp_francke_ree2024. The Lifestyle spread captures amenity-driven cycles in coastal and regional destinations, where markets often lag the Market baseline for extended periods and then catch up sharply during migration-driven upswings.
This paper contributes a transparent three-factor decomposition in levels that is tightly linked to PCA evidence yet constructed from simple index spreads. Because the factors are portable—reproducible from standard city/region price indexes supplied by multiple vendors. Estimating the model via ARIMAX, we show that these three factors capture most of the low-frequency co-movement and the dominant deviations in Australian city house prices over the last 30 years.
@@
This stability enables transparent scenario mapping: for a multiplicative change $f_{M}$ in the national price index, the market-only regional change is approximately $f_{M}^{\beta_{r}}$, with uncertainty quantified by factor-specific forecast funnels (edwards2007). In the data, Sydney and Canberra (ACT) track the national market closely ($\beta_{r}\approx 1$), Melbourne and Brisbane amplify national movements, and regional areas dampen them. We emphasize that these relationships are conditional on the national path: while the national Market trend itself may evolve, the relative scaling via $\beta_{r}$ and the mean-reverting character of the spreads show no evidence of trend instability in expanding windows over 1995–2024, suggesting that the estimated loadings and forecast uncertainties are likely to remain informative in the near future of 5 to 10 years.
@@
A long-standing user-cost literature studies the level of house prices relative to rents, interest rates and taxes, typically within an asset-market framework (poterba1984; poterba1992; himmelberg_mayer_sinai2005; campbell_etal2009). This body of work highlights the role of real mortgage rates, expected capital gains and tax treatment in shaping housing valuations and the rent–price ratio. In the Australian context, empirical studies have related house prices to fundamentals such as income, interest rates and supply constraints (e.g. abelson1994; abelson_etal2005; bodman_crosby2004; otto2007; fox_tulip2014; saunders_tulip2019; abelson_joyeux2023). Consistent with wilson_zurbruegg2008, who use cointegration and restriction tests to show that the long-run fundamentals driving house prices differ across Australian capitals, we explicitly examine how long-run growth and factor exposures vary by city. Our focus and method are different, however: rather than modelling user cost or long-run equilibrium prices as functions of explanatory variables, we take the house price data as a starting point and analyse how a small number of common and geographically structured factors account for differences in long-run price growth across cities.
otto2007 documents substantial dispersion in real house price growth and volatility across Australian capitals over 1986–2005, using ABS indexes for established houses. Our study differs in both period and methodology: we analyze a longer and more recent window (1995–2024), construct repeat-sales indexes from transaction data, and employ a factor model that decomposes a national Market trend from Mining and Lifestyle spreads. He also confirms earlier work by abelson1994 that Sydney prices tend to lead some of the other cities, which we interpret as evidence of common shocks. In our framework these are captured by a symmetric national Market factor and two geographically structured spread factors (Mining and Lifestyle), so Sydney is just one market loading on shared influences rather than an exogenous driver.
@@
In examining the last U.S. housing boom, delnegro2007 employ a Bayesian dynamic factor model in the spirit of geweke1977 on quarterly OFHEO house price indices for U.S. states (1986–2005) to disentangle the common component from the region-specific shocks, finding an increasingly concerted (upward) price movement towards the end of their period (2001–2004), an important period containing the U.S. housing boom of the early 2000s, indicating more heterogeneous behavior in earlier periods. Their returns-based framework isolates a common (national) housing factor from local (state- or region-specific) shocks using fixed $\beta$ coefficients for the cities. This establishes two ideas we adopt: first, using a national housing factor to capture broad comovement across cities; and second, using a small set of factors to capture broader geographically varying growth similarities. In contrast to their model, we do not employ a dynamic equation of motion for these factors, but rather base them on time series that are known to explain most of the dataset variance.
@@
case2010housing adopt a risk-related asset-pricing framework as in finance, modeling quarterly MSA returns against a (differenced) national housing “market” factor (proxied by the aggregate U.S. house price index) and standard risk characteristics (size/SMB, momentum, and idiosyncratic risk). They find a sizable market effect and, using Fama–MacBeth portfolio tests, evidence that idiosyncratic (non-market) risk also helps price housing returns, consistent with a positive risk–return trade-off view a la fama_french1992. A closely related perspective in the Australian context is offered by morawakage2022, who, after accounting for spatial dependence and heterogeneity in the Brisbane market, report that at this more granular level, idiosyncratic (asset-specific/locational) risks explain excess returns at the submarket (postcode) level, with common speculative forces also playing a role.  [^2] For a broader review of return-based housing asset pricing, see morawakage2022 and references therein. Our focus is on levels-based growth decomposition rather than the pricing of risk in returns.
@@
In contrast to these risk-related asset-pricing models formulated in returns, our objective is to characterize persistent relative growth differentials over multi-decadal periods at the aggregated city level, not the cross-sectional pricing of risk. We retain a national market component but operate in levels to preserve long-run co-movement (cf. holly_etal2010) and, rather than risk-priced characteristics, we use geographically oriented factors anchored in the PCA evidence of sijp_francke_ree2024. Consistent with user-cost perspectives that view owner-occupied housing primarily as a provider of services rather than a traded financial asset (himmelberg_mayer_sinai2005), our factor model is phenomenological: it isolates stable long-run growth sensitivities and mean-reverting geographic cycles without claiming a structural mechanism for their origin.
@@ 2 Methods and Methodology
In this paper, we introduce a three-factor model for price indexes of regions within Australia. The model includes a national market trend that captures the main growth trajectory and links local price dynamics to national developments, along with two mean-reverting components. This section outlines the model’s construction and data used in its empirical implementation.
@@ 2.1 Data
As a brief aside, while we report city-level results for interpretability, Sydney exhibits within-city variation. Using SA2 indices (Jan 1995 = 1), the median SA2 has grown $8.83\times$; the interquartile range is $8.10\times,\,9.86\times]$, with the central 80% spanning $7.48\times,\,11.15\times]$. A small right tail (4.3% of SA2s) exceeds $12\times$, plausibly reflecting localized shocks (such as infrastructure, rezoning, or redevelopment). These pockets widen forecast bands through the idiosyncratic remainder $\epsilon_{r}$, but do not overturn $\beta_{r}$ rankings or our aggregate conclusions.
@@ 2.2 Approximating price indexes using PCA
sijp_francke_ree2024 apply PCA to Australian regional house price indexes and identify three dominant components. We replicate their analysis on the SA2 price-index collection described in Section Appendix B and recover the same three leading components, which we label Market (PC1), Mining (PC2), and Lifestyle (PC3); their economic interpretation is summarised in Section 2.3. We verify that the first component $z_{1}$ tracks the national price index $U$ (the weighted mean of SA2 indexes) almost identically, with correlation exceeding 0.99 (Table 2).
As PCA is linear, any regional log price index $\mu_{r}$ can be approximated by a linear combination of principal components:
@@ 2.3 Economic interpretation of the leading PCs
PC1 (“Market”) is a broad national component capturing the common drift in dwelling values across population centres. Regions with higher loadings on PC1 tend to move more than one-for-one with aggregate Australian housing cycles.
PC2 (“Mining”) reflects resource-sector dynamics. Correlations with mining investment indicate that PC2 is most strongly expressed in resource-exposed regions (e.g. WA, QLD). Its time profile tracks the 2000s mining-investment boom and its aftermath, consistent with major macro shocks over that period (downes_et_al_rba_report2014). Labor migration during the mining boom (foo_salim_minecon2022) likely contributed to these regional price pressures. Loadings on this component are negative in Sydney (and to a lesser extent Melbourne), indicating that, relative to mining-exposed regions, these markets exhibit weaker or opposite-signed comovement with the resource cycle. See Section 3.4 for further discussion.
PC3 (“Lifestyle”) loads positively on amenity-rich coastal and regional areas associated with “sea/tree-change” migration rather than employment centres. This reflects a search for a more affordable lifestyle, often in scenic locations, with a better work-life balance, increasingly enabled by remote work. Evidence documents sustained household shifts from capitals toward regional lifestyle destinations, intensifying during and after COVID-19 (bohnet2010; yanotti2023), with spillovers from inner capital-city markets into the regions. Consistent with this, lifestyle regions move together in long cycles: they lag capitals during booms, then catch up rapidly near the end, experiencing affordability pressure as migration inflows rise (sijp2025report). Listing-text evidence supports this link: high-PC3 regions show elevated frequencies of lifestyle descriptors (e.g., beach, coastal, lifestyle). [^3] See sijp2025report for the word-count study linking listing text to PC3. This internal migration phenomenon also agrees with the common knowledge of the long-standing pattern of households moving from the colder southern states to the warmer state of Queensland and coastal regions of NSW. Our data indicates a wide and persistent influence within Australia. Note that these loadings are produced by the PCA, they are not a-priori identified with mining and lifestyle.
@@ 2.4 Identifying the factors for the factor model
The linear approximation of regional indexes by PCA components (Equation 1) motivates a factor model in which each $\mu_{r}$ is expressed as a linear combination of a small number of interpretable factors plus an autocorrelated residual. Our aim is to construct such a model using simple, readily available factors derived from aggregate price indexes that proxy the leading principal components while remaining stable under expanding estimation windows. We instead seek concrete differences in broad indexes between locations that replicate the PC behavior but are grounded in intuitive price differentials and are portable: the Market index and Mining/Lifestyle differences can be reproduced directly from standard city or regional price indexes available from multiple data vendors, making the framework easy to implement on other datasets without the need for a large underlying national collection of granular indexes. We refer to the Mining and Lifestyle factors as spreads, in the sense that they capture mean-reverting differentials in regional prices relative to the national Market trend.
@@
PC1, $z_{1}$, is nearly identical to the national price index $U$, we therefore adopt $U$ as our Market factor. This represents the common trend across all regions and absorbs the unit-root (I(1)) component of the system. Here, we have used the PCA procedure to guide our choice of market factor, and the PCA is independent of the weights used to construct a national index: it is solely based on the co-movement of the regions irrespective of weight.
@@
yielding $\alpha=0.94$ for our data. This ratio ensures that the spread $\delta_{PS}$ has a low correlation with the Market factor $U$, isolating the Perth-Sydney differential net of common national growth. Perth’s economy is heavily resource-dependent while Sydney’s is services-dominated and has the most negative PCA 2 loadings, rendering their relative performance a natural proxy for resource-sector cycles. Figure 1a compares $\delta_{PS}$ to PC2; the two series track each other closely, with a correlation of 0.98 (Table 2).
@@
For PC3, which sijp_francke_ree2024 associate with lifestyle and amenity-driven demand, there is no single city-level spread that naturally proxies the component. Instead, we construct a Lifestyle spread by taking an average weighted by the number of houses (using the weights defined in Appendix B) of the top 20 lifestyle SA4 regions according to their PC3 loadings, and offsetting this against the bottom 20 regions. [^4] Top Lifestyle regions (positive PC3 loadings): Central Coast, Hunter Valley Exc Newcastle, South Australia - South East, Ipswich, Capital Region, Hobart, Central West, South East, Moreton Bay - North, Riverina, Southern Highlands And Shoalhaven, Richmond - Tweed, Coffs Harbour - Grafton, Murray, Gold Coast, West And North West, Launceston And North East, Sunshine Coast, Wide Bay, Mid North Coast. Bottom Lifestyle regions (negative PC3 loadings): Sydney Inner West, Western Australia - Wheat Belt, Sydney City And Inner South, Perth - Inner, Melbourne - Outer East, Perth - North West, Melbourne - South East, Western Australia - Outback (south), Perth - South East, Perth - North East, Melbourne - North West, Melbourne - North East, Melbourne - West, Melbourne - Inner South, Melbourne - Inner East, Melbourne - Inner, Western Australia - Outback (north), Queensland - Outback, Darwin, Northern Territory - Outback. As with the Mining spread, we apply a trend-adjustment coefficient $\alpha=0.82$ to remove the common market factor. Figure 1b shows the Lifestyle spread alongside PC3; the two series track closely, with a correlation of 0.98 (Table 2).
@@
Finally, Table 3 shows that the three factors are nearly orthogonal, with pairwise correlations between -0.18 and 0.19. This near-independence simplifies estimation and interpretation: each factor contributes additively to regional price dynamics without strong multicollinearity. Note that the trend‐adjustment coefficients $\alpha$ were chosen based on criteria independent to the model selection presented in later sections, and were held fixed throughout all specifications.

### v3

@@
December 14, 2025
Australian house prices have risen strongly since the mid-1990s, but growth has been highly uneven across regions. Raw growth figures obscure whether these differences reflect persistent structural trends or cyclical fluctuations. We address this by estimating a three-factor model in levels for regional repeat-sales log price indexes over 1995–2024. The model decomposes each regional index into a national Market factor, two mean-reverting factors (Mining and Lifestyle) that capture decadal cycles that are expressed differently geographically, and a city-specific residual. The Mining factor, proxied by a Perth–Sydney index differential, reflects resource-driven cycles in relative performance; the Lifestyle spread captures amenity-driven coastal and regional pull and the associated city push. The Market loading $\beta_{r}$ isolates each region’s fundamental sensitivity to national growth, so that an assumed national change $f_{M}$ implies an individual city’s growth of $f_{r}\approx f_{M}^{\beta_{r}}$ once mean-reverting spreads are netted out. Comparing realised paths to these factor-implied trajectories indicates when a city is historically elevated or depressed, and attributes the gap to Mining or Lifestyle spreads.
Expanding-window ARIMAX estimation reveals that Market betas are stable across major shocks (the mining boom, the Global Financial Crisis, and COVID-19), while Mining and Lifestyle behave as stationary spreads that widen forecast fans without overturning the cross-sectional ranking implied by $\beta_{r}$. Once corrected for the geographic cycles, Melbourne amplifies national growth, Sydney tracks the national trend closely, and regional areas dampen it. The framework thus provides a simple, factor-based tool for interpreting regional growth differentials and their persistence.
@@ 1 Introduction
Australian house price appreciation has sustained considerable variation across the capital cities and the regional areas over decades. Local housing markets are generally shaped by region-specific conditions, while sharing a common national price trend. This paper studies such divergence for Australian houses (excluding units and apartments) over 1995–2024, using nominal repeat-sales log price indexes. We examine how much of the long-run divergence across cities reflects broadly shared but geographically varying cyclical forces, and how much instead reflects each city’s long-run sensitivity to the national market trend. These regional growth differentials matter for housing policy, investment decisions, and spatial inequality: persistent gaps can impede internal migration and exacerbate local skill shortages, while reliable planning requires understanding whether relative growth relationships are stable.
Our approach takes the PCA study of sijp_francke_ree2024 as a starting point, and separates a trend correlated to the national price index, the Market factor, from two economically interpretable, mean-reverting factors that are expressed differently at various locations. The Mining spread captures the geographic variation in exposure to the resource cycle that raises prices more in mining-exposed regions, relative to the national market and less in Sydney, consistent with its negative Mining factor exposure in both our factor model and the PCA of sijp_francke_ree2024. The Lifestyle spread captures amenity-driven cycles in coastal and regional destinations, where markets often lag the Market baseline for extended periods and then catch up sharply during internal migration-driven upswings.
This paper contributes a transparent three-factor decomposition in levels that is tightly linked to PCA evidence, yet constructed from simple (weighted) index differences, or “spreads”. The factors are “portable” in the sense that they are reproducible from standard city/region price indexes supplied by multiple vendors. Estimating the model via ARIMAX, we show that these three factors capture most of the low-frequency co-movement and the dominant deviations in Australian city house prices over the last 30 years.
@@
This stability enables transparent scenario mapping: for a multiplicative change $f_{M}$ in the national price index, the market-only regional change is approximately $f_{M}^{\beta_{r}}$, with uncertainty quantified by factor-specific forecast fans (edwards2007). In the data, Sydney and Canberra (ACT) track the national market closely ($\beta_{r}\approx 1$), Melbourne and Brisbane amplify national movements, and regional areas dampen them. We emphasize that we do not predict the national path (or even the remaining factors). Instead, we examine city price developments relative to the national index. While the national Market trend itself may evolve, the relative scaling via $\beta_{r}$ and the mean-reverting character of the spreads show no evidence of trend instability in expanding windows over 1995–2024, suggesting that the estimated loadings and forecast uncertainties are likely to remain informative in the near future of 5 to 10 years. We do not attempt to predict the national path. Any reference to national growth is a descriptive background rather than a forward-looking projection.
@@
We briefly outline some of the literature on variations in house price growth and underlying drivers. Price trajectories can diverge substantially, as evidenced by asynchronous booms across U.S. cities during the 2000s housing cycle (ferreira_gyourko2011). Individual city growth can exceed the national average over very long periods of time (gyourko_mayer_sinai2013). As to the drivers, a long-standing user-cost literature studies the level of house prices relative to rents, interest rates and taxes, typically within an asset-market framework (poterba1984; poterba1992; himmelberg_mayer_sinai2005; campbell_etal2009). This body of work highlights the role of real mortgage rates, expected capital gains and tax treatment in shaping housing valuations and the rent–price ratio. In the Australian context, empirical studies have related house prices to fundamentals such as income, interest rates and supply constraints (e.g. abelson1994; abelson_etal2005; bodman_crosby2004; otto2007; fox_tulip2014; saunders_tulip2019; abelson_joyeux2023). wilson_zurbruegg2008 use cointegration and restriction tests to show that the long-run fundamentals driving house prices differ across Australian capitals. otto2007 documents substantial dispersion in real house price growth and volatility across Australian capitals over 1986–2005, using ABS indexes for established houses. He also confirms earlier work by abelson1994 that Sydney prices tend to lead some of the other cities, which we interpret as evidence of common shocks. In our framework these are captured by a symmetric national Market factor and two geographically structured spread factors (Mining and Lifestyle), so Sydney is just one market loading on shared influences rather than an exogenous driver. We leave the explicit accommodation of leading cities and spillovers within our factor framework for future work. More generally, our focus and approach differs from the user-cost related literature. Rather than modelling user cost or long-run equilibrium prices as functions of explanatory variables, we take the house price data as a starting point and analyse how a small number of common and geographically structured factors account for differences in long-run price growth across cities.
@@
In examining the last U.S. housing boom, delnegro2007 employ a Bayesian dynamic factor model in the spirit of geweke1977 on quarterly OFHEO house price indices for U.S. states (1986–2005) to disentangle the common component from the region-specific shocks, finding an increasingly concerted (upward) price movement towards the end of their period (2001–2004). This is an important period as it encompasses the U.S. housing boom of the early 2000s. Their returns-based framework isolates a common (national) housing factor from local (state- or region-specific) shocks using fixed $\beta$ coefficients for the cities. This establishes two ideas we adopt: first, using a national housing factor to capture broad comovement across cities; and second, using a small set of factors to capture broader geographically varying growth similarities. In contrast to their model, we do not employ a dynamic equation of motion for these factors, but rather base them on time series that are known to explain most of the dataset variance and are associated with broad socio-economic phenomena.
@@
case2010housing adopt a risk-related asset-pricing framework, inspired by such models in finance, and model quarterly MSA returns against a (differenced) national housing “market” factor (proxied by the aggregate U.S. house price index), and standard risk characteristics (size/SMB, momentum, and idiosyncratic risk). They find a sizable market effect and, using Fama–MacBeth portfolio tests, evidence that idiosyncratic (non-market) risk also helps price housing returns, consistent with a positive risk–return trade-off view a la fama_french1992. A related perspective in the Australian context is offered by morawakage2022, who, after accounting for spatial dependence and heterogeneity in the Brisbane market, report that at this more granular level, idiosyncratic (asset-specific/locational) risks explain excess returns at the submarket (postcode) level, with common speculative forces also playing a role. [^2] For a broader review of return-based housing asset pricing, see morawakage2022 and references therein. Our focus is on levels-based growth decomposition rather than the pricing of risk in returns.
@@
In contrast to these risk-related asset-pricing models formulated in returns, our objective is to characterize persistent relative growth differentials over multi-decadal periods at the aggregated city level, not the cross-sectional pricing of risk. We retain a national market component but operate in levels to preserve long-run co-movement (cf. holly_etal2010) and, rather than risk-priced characteristics, we use geographically oriented factors anchored in the PCA evidence of sijp_francke_ree2024. Without taking a traded financial asset view (case2010housing) or an alternative user-cost perspective that views owner-occupied housing primarily as a provider of services (himmelberg_mayer_sinai2005), our factor model is descriptive and remains compatible with either philosophy.
@@ 2 Methods and Methodology
This section outlines the construction of the three-factor model for price indexes of regions within Australia and data used in its empirical implementation.
@@ 2.1 Data
As a brief aside, while we report city-level results for interpretability, Sydney exhibits within-city variation. Using SA2 indices (Jan 1995 = 1), the median SA2 has grown $8.83\times$; the interquartile range is $8.10\times,\,9.86\times]$, with the central 80% spanning $7.48\times,\,11.15\times]$. A small right tail (4.3% of SA2s) exceeds $12\times$, plausibly reflecting localized shocks (such as infrastructure, rezoning, or redevelopment). These pockets widen the forecast bands through the idiosyncratic remainder $\epsilon_{r}$, but do not overturn $\beta_{r}$ rankings or our aggregate conclusions.
@@ 2.2 Approximating price indexes using PCA
sijp_francke_ree2024 apply PCA to Australian regional house price indexes and link the three most dominant components to specific socio-economic processes. We replicate their analysis on the SA2 price-index collection described in Section Appendix B and recover the same three leading components, which we label Market (PC1), Mining (PC2), and Lifestyle (PC3); their economic interpretation is summarised in Section 2.3.
As PCA is linear, any regional log price index $\mu_{r}$ can be approximated by a linear combination of leading principal components:
@@ 2.3 Economic interpretation of the leading PCs
PC1 (“Market”) is a broad national component capturing the common drift in dwelling values across population centres. Regions with higher loadings on PC1 tend to move more than one-for-one with aggregate Australian housing cycles. We verify that the first component $z_{1}$ tracks the national price index $U$ (the weighted mean of SA2 indexes) almost identically, with correlation exceeding 0.99 (Table 2).
PC2 (“Mining”) reflects resource-sector dynamics. Correlations with mining investment (calculated in sijp_francke_ree2024) indicate that PC2 is most strongly expressed in resource-exposed regions (e.g. WA, QLD). Its time profile tracks the 2000s mining-investment boom and its aftermath, consistent with major macro shocks over that period (downes_et_al_rba_report2014). Labor migration during the mining boom (foo_salim_minecon2022) likely contributed to these regional price pressures. Loadings on this component are negative in Sydney (and to a lesser extent Melbourne), indicating that, relative to mining-exposed regions, these markets exhibit weaker or opposite-signed comovement with the resource cycle. See Section 3.4 for further discussion.
PC3 (“Lifestyle”) loads positively on amenity-rich coastal and regional areas associated with “sea/tree-change” migration rather than employment centres. This reflects a search for a more affordable lifestyle, often in scenic locations, with a better work-life balance, increasingly enabled by remote work. Evidence more generally documents households moving from capital cities toward regional towns, including lifestyle destinations (bohnet2010; yanotti2023). This has intensified during and after COVID-19, but the pattern is evident for Lifestyle regions throughout the last 30 years, with the penultimate upswing in 2000-2004. These regions lag capitals during booms, then catch up rapidly near the end, experiencing affordability pressure as migration inflows (from the cities) rise (sijp2025report). This movement also agrees with the common knowledge of the long-standing pattern of households moving from the colder southern states to the warmer state of Queensland and coastal regions of NSW. Among the evidence to link PC3 to lifestyle is the correlation between PC3 loadings and word frequencies in real estate advertisement listings: high-PC3 regions show elevated frequencies of lifestyle descriptors (e.g., beach, coastal, lifestyle). [^3] See sijp2025report for the word-count study linking listing text to PC3. Note that these loadings are produced by the PCA, they are not a-priori identified with mining and lifestyle.
@@ 2.4 Identifying the factors for the factor model
The linear approximation of regional indexes by PCA components (Equation 1) motivates a factor model in which each $\mu_{r}$ is expressed as a linear combination of a small number of interpretable factors plus an autocorrelated residual. Our aim is to construct such a model using simple, readily available factors derived from aggregate price indexes that proxy the leading principal components while remaining stable under expanding estimation windows. To this end, we construct the non-trend factors in terms of concrete differences in broad indexes between locations that replicate the PC behavior but are grounded in intuitive price differentials and are portable: the Market index and Mining/Lifestyle differences can be reproduced directly from standard city or regional price indexes available from multiple data vendors, making the framework easy to implement on other datasets without the need for a large underlying national collection of granular indexes.
For clarity, we observe that the PCA derived series $z_{1}$, $z_{2}$ and $z_{3}$ correspond directly to the first three principle components: PC1 (Market), PC2 (Mining), PC3 (Lifestyle) respectively. The factor proxies introduced below are constructed explicitly to serve as empirically stable approximations to these components.
@@
PC1, $z_{1}$, is nearly identical to the national price index $U$, we therefore adopt $U$ as our Market factor. This represents the common trend across all regions and absorbs the unit-root component of the system. Here, we have used the PCA procedure to guide our choice of market factor, and the PCA is independent of the weights used to construct a national index: it is solely based on the co-movement of the regions irrespective of weight.
@@
yielding $\alpha=0.94$ for our data. This ratio ensures that the spread $\delta_{PS}$ has a low correlation with the Market factor $U$, isolating the Perth-Sydney differential net of common national growth. Among the candidate bilateral spreads evaluated, the Perth–Sydney differential demonstrated the strongest empirical correspondence with PC2. Perth’s economy is heavily resource-dependent while Sydney’s is services-dominated and has the most negative PCA 2 loadings, rendering their relative performance a natural proxy for resource-sector cycles. Figure 1a compares $\delta_{PS}$ to PC2; the two series track each other closely, with a correlation of 0.98 (Table 2).
@@
For PC3, there is no single city-level spread that naturally proxies the component. Instead, we construct a Lifestyle spread by taking an average weighted by the number of houses (using the weights defined in Appendix B) of the top 20 lifestyle SA4 regions according to their PC3 loadings, and offsetting this against the bottom 20 regions. This isolates the strongest lifestyle signal while avoiding noise from weakly loaded regions.  [^4] Top Lifestyle regions (positive PC3 loadings): Central Coast, Hunter Valley Exc Newcastle, South Australia - South East, Ipswich, Capital Region, Hobart, Central West, South East, Moreton Bay - North, Riverina, Southern Highlands And Shoalhaven, Richmond - Tweed, Coffs Harbour - Grafton, Murray, Gold Coast, West And North West, Launceston And North East, Sunshine Coast, Wide Bay, Mid North Coast. Bottom Lifestyle regions (negative PC3 loadings): Sydney Inner West, Western Australia - Wheat Belt, Sydney City And Inner South, Perth - Inner, Melbourne - Outer East, Perth - North West, Melbourne - South East, Western Australia - Outback (south), Perth - South East, Perth - North East, Melbourne - North West, Melbourne - North East, Melbourne - West, Melbourne - Inner South, Melbourne - Inner East, Melbourne - Inner, Western Australia - Outback (north), Queensland - Outback, Darwin, Northern Territory - Outback. As with the Mining spread, we apply a trend-adjustment coefficient $\alpha=0.82$ to remove the common market factor. Figure 1b shows the Lifestyle spread alongside PC3; the two series track closely, with a correlation of 0.98 (Table 2).
@@
Finally, Table 3 shows that the three factors are nearly orthogonal, with pairwise correlations between -0.18 and 0.19. This near-independence simplifies estimation and interpretation by ensuring that each factor contributes additively to regional price dynamics without strong multicollinearity, and remains consistent with the orthogonality of the underlying principal components. Note that the trend‐adjustment coefficients $\alpha$ were chosen based on criteria independent to the model selection presented in later sections, and were held fixed throughout all specifications.